# Unifying Summary Statistic Selection for Approximate Bayesian Computation

### v3

@@ 4.1 Approximate sufficiency
Joyce & Marjoram (2008) cast the task of selecting summaries as a sequence of hypothesis tests to select a subset of candidate summaries. Specifically, they considered
$\log R_{k}\left(\theta\right)=\log f\left(\theta\mid t_{k},\ldots,t_{1}\right)-\log f\left(\theta\mid t_{k-1},\ldots,t_{1}\right),$
@@
where $f\left(\theta\mid t_{k-1},\ldots,t_{1}\right)$ is the approximate posterior given $k-1$ summaries already selected and $f\left(\theta\mid t_{k},\ldots,t_{1}\right)$ is the posterior resulting from including an additional statistic $t_{k}$. Intuitively, if the error score $\delta_{k}=\max_{\theta}\left|\log R_{k}\left(\theta\right)\right|$ is zero, i.e. the two posteriors are identical, the $k^{\mathrm{th}}$ statistic does not capture additional information and can be ignored. If $\delta_{k}$ differs significantly from zero, we reject the null hypothesis that $f\left(\theta\mid t_{k},\ldots,t_{1}\right)$ and $\log f\left(\theta\mid t_{k-1},\ldots,t_{1}\right)$ are the same distribution and include $t_{k}$.
@@
However, the error score $\delta_{k}=\max_{\theta}\left|\log R_{k}\right|$ assigns equal importance to all subsets of the parameter space, even regions we know to be irrelevant. For example, suppose that the posterior given the currently selected $k-1$ summaries is normal with variance $\sigma_{k-1}^{2}$, and the posterior after adding the $k^{\mathrm{th}}$ summary is identical except for a different variance $\sigma_{k}^{2}$. Even if $\sigma_{k}$ and $\sigma_{k-1}$ differ by an infinitesimal amount, $\delta_{k}$ is unbounded because
@@
$\delta_{k}=\frac{1}{2}\max_{\theta}\left|\log\left(\frac{\sigma_{k}^{2}}{\sigma_{k-1}^{2}}\right)+\left(\frac{\sigma_{k-1}^{2}-\sigma_{k}^{2}}{\sigma_{k}^{2}\sigma_{k-1}^{2}}\right)\theta^{2}\right|=\infty.$
@@
The error score is dominated by regions of the parameter space that have virtually no posterior mass. The expected value $\mathbb{E}_{\theta\sim f\left(\theta\mid t_{k},\ldots,t_{1}\right)}\left[\log R_{k}\left(\theta\right)\right]$ instead weights discrepancies between the two distributions by the posterior mass. This quantity is in fact the KL divergence considered by Barnes et al. (2012) (see Section 4.5 for details).
@@
The notion of “approximate” sufficiency is necessarily a statement about limited computational resources: If we had unlimited resources, only candidate statistics that are uninformative or redundant would be excluded. This observation applies to any subset selection algorithm, such as minimizing posterior entropy (Nunes & Balding, 2010) in Section 4.2, regression-based subset selection methods (Blum & François, 2010; Blum et al., 2013) in Section 4.4, or maximizing MI (Barnes et al., 2012) in Section 4.5.
@@ 4.2 Minimizing the conditional posterior entropy
Nunes & Balding (2010) proposed choosing a subset of summaries $t$ by minimizing the conditional posterior entropy (CPE) $H\left\{f\left(\theta\mid t\left(y\right)\right)\right\}$ given data $y$. They ran rejection ABC for different subsets of summaries and evaluated the CPE using a nearest-neighbour estimator (Singh et al., 2003). The proposal is appealing because low-entropy posteriors give precise parameter estimates.
However, it implicitly assumes that the data we have observed are the only data that could ever be observed, similar to the non-parametric bootstrap. When the maximum likelihood estimate of the parameters lies in the tail of the prior distribution, the CPE $H\left\{f\left(\theta\mid y\right)\right\}$ can be larger than the prior entropy $H\left\{\pi\left(\theta\right)\right\}$ because the posterior is a “compromise” between prior and likelihood (Blum et al., 2013).
@@
[tough-nuts]
Figure 2: Extracting summaries can be non-trivial even for toy models. Panel (a) shows the difference between posterior and prior entropy for a model with zero-mean normal likelihood and conjugate gamma prior for the precision $\theta$ (inverse variance). For a subset of the prior and data space, minimizing the posterior entropy discards the second moment $t$, a sufficient statistic. Panel (b) shows the bimodal posterior for the example point in (a) that arises when the precision of the likelihood is $\left|\theta\right|$ (see Section 4.4). The posterior mean is zero and not informative of the parameter. The vertical dashed line represents the maximum likelihood estimate $\widehat{\left|\theta\right|}$ of the precision $\left|\theta\right|$.
@@
where $t=n^{-1}\sum_{i=1}^{n}y_{i}^{2}$ is the second moment, a sufficient statistic. For example, if $a=1.5$ and $t=0.3$, the prior entropy is $1.36$ and the CPE is $1.47$. Minimizing the CPE would discard the sufficient statistic $t$ such that the posterior is equal to the prior: We have not learned anything from the data. Panel (a) of Fig. 2 shows the entropy gain $\Delta=H\left\{f\left(\theta\mid y\right)\right\}-H\left\{\pi\left(\theta\right)\right\}$ in light of the data for different priors and sample variances. Indeed, generating data from the prior predictive distribution with $a=1.5$, we find that $30\%$ of samples lead to a CPE increase. Interestingly, this situation is more likely to arise when the “surprise” (Itti & Baldi, 2009) is large, and we should substantially update our beliefs in light of the data. In contrast, the EPE $\mathcal{H}=0.87$ given $t$ is smaller than the prior entropy, and minimizing it would select $t$ as a useful summary. We estimated the EPE and proportion of entropy increases by drawing $10^{5}$ samples from the generative model. Monte Carlo standard errors are smaller than the reported significant digits.
Nunes & Balding (2010) also considered a two-stage method: First they used the above approach to select candidate summaries and identify simulated datasets close to the observed data. Second, they drew posterior samples for each identified dataset and evaluated the root mean integrated squared error (RMISE) of posterior samples for each subset of summaries. This is possible because the parameters of simulated datasets are known. The summaries with the lowest RMISE were then selected. We do not consider this two-stage approach further here because of its computational burden and because posterior mean estimation methods optimize a similar objective, as discussed in Section 4.4.
@@ 4.3 Maximizing the Fisher information
Even when the likelihood is tractable, compressing the data $y$ to summaries $t$ has computational benefits. Heavens et al. (2000) developed an optimal linear compression scheme for Gaussian likelihoods in the sense that the Fisher information is preserved. Information-maximizing neural networks (Charnock et al., 2018) seek to maximize the Fisher information when linear compression is not sufficient, and methods to maximize the Fisher information for non-Gaussian likelihoods have recently been developed (Alsing & Wandelt, 2018; Fluri et al., 2021).
We consider the large-sample limit $n\rightarrow\infty$ such that, according to the Bernstein–von Mises theorem, the posterior approaches a multivariate normal distribution under certain regularity conditions which we will consider further (van der Vaart, 1998). Specifically,
@@
is the Fisher information of the summaries evaluated at $\theta_{0}$ (Bishop, 2006, Ch. 6). The limiting entropy of the posterior can thus be readily evaluated and is
@@
We argue that minimizing the EPE is more appealing than maximizing the Fisher information for three reasons. First, it can incorporate prior information in the small-$n$ regime to yield the most faithful posterior approximation. Second, it does not require the choice of a fiducial value of $\theta$ at which to evaluate the Fisher information. Finally, when the likelihood is not available, we need to approximate it to evaluate the Fisher information. For example, Charnock et al. (2018) assume that the likelihood of the learned summaries can be approximated by a Gaussian, and Alsing & Wandelt (2018) argue that candidate summaries often have a Gaussian likelihood if they are the mean of i.i.d. data.
@@ 4.4 Minimizing the Bayes risk
Fearnhead & Prangle (2012) proposed the posterior mean of the parameters as summaries. Of course, the posterior mean is not known, but we can estimate it by minimizing the quadratic loss
@@
where $t_{\beta}(z)$ is a predictor of $\theta$ parameterized by $\beta$, $A$ is a positive-definite matrix, and  denotes the transpose. Fearnhead & Prangle (2012) considered linear predictors, but neural networks (Jiang et al., 2017) and boosted regression (Aeschbacher et al., 2012) have also been proposed. In practice, the parameters $\beta$ are learned by minimizing a Monte Carlo estimate of Eq. 9 akin to Eq. 4. Using the estimated posterior mean $t_{\beta}\left(\cdot\right)$ as summaries implicitly chooses as many summaries as there are parameters.
Considering again the large-sample limit, the quadratic loss becomes (adapted from Theorem 3 of Fearnhead & Prangle (2012))
@@
The above argument crucially depends on the assumptions of the Berstein–von Mises theorem holding. In particular, the model needs to be identifiable such that different values of the parameters $\theta$ are distinguishable in the $n\rightarrow\infty$ limit (van der Vaart, 1998). We consider a variant of the toy model presented in Section 4.2 that is not identifiable and discuss the impact on learning summaries. In particular, we use the absolute value $\left|\theta\right|$ of a parameter $\theta$ as the precision such that the conditional distributions are
$\theta\mid a,b$ $\sim u\times\mathsf{Gamma}\left(a,b\right)$
$y_{i}$ $\sim\mathsf{Normal}\left(0,\left|\theta\right|^{-1}\right),$
where $u$ is ±$\pm 1$ with equal probability. The posterior is
$\theta\mid y,a,b\sim u\times\mathsf{Gamma}\left(a+\frac{n}{2},b+\frac{nt}{2}\right),$
@@
This example may seem contrived, but multimodal posteriors that render the posterior mean uninformative are not uncommon. For example, mixture models are invariant under label permutation (Stephens, 2000), and latent-space models of networks (Hoff et al., 2002) as well as latent factor models for Bayesian PCA (Nirwan & Bertschinger, 2019) are invariant under rotations. Using information theoretic approaches ensures we stay focused on the task at hand: Approximating the posterior.
The relationship between parameters and data can be complex, and regression approaches, especially linear regression, may not be able to capture the relationship globally. Local relationships in regions of high posterior mass can be learned using pilot runs (Fearnhead & Prangle, 2012) or weighting samples (Blum & François, 2010). Local regression methods have also been adapted for subset selection: A model is fit to predict parameters from candidate summaries, and a candidate is selected if it increases a metric such as the Bayesian evidence (Blum & François, 2010), Akaike information criterion, or Bayesian information criterion (Blum et al., 2013).
@@ 4.5 Maximizing the mutual information
Barnes et al. (2012) proposed choosing summaries from a pool of candidates that maximize the MI $I\left\{\theta,t\right\}$ between parameters $\theta$ and the statistics $t$. Assuming that the candidate set includes sufficient statistics $t_{\text{suff}}$ such that
@@
for all possible $y$, they constructed a set of summaries sequentially. At the $k^{\mathrm{th}}$ step, they included the summary that maximizes the surprise given the $k-1$ statistics that have already been selected. The approach is similar to the approximate sufficiency method reviewed in Section 4.1, but candidates are prioritized by their surprise at each stage. Together, the steps select the summaries that maximize the surprise $D_{\mathrm{KL}}\left(f\left(\theta\mid t\right)\;\|\;\pi\left(\theta\right)\right)$.
However, recall from Eqs. 5 and 6 that the MI is equal to the expected surprise under the generative model. In general, maximizing the surprise for a particular observed dataset is thus not equivalent to maximizing the MI. The approach may select different summaries if the candidate set does not include sufficient statistics.
Similarly, Chen et al. (2021) sought to maximize the MI using a neural network. They suggested that “$t(z)$ is a sufficient statistic for $g\left(z\mid\theta\right)$ if and only if” (p. 2) it maximizes the MI and “that the sufficiency of the learned statistics is insensitive to the choice of $\pi\left(\theta\right)$” (p. 4) such that “[their approach] is globally sufficient for all $\theta$” (p. 6) [^1] We have adapted notation in quotations for consistency with this manuscript.. As we shall illustrate with a toy model, these propositions do not hold in general because of the difference between sufficient and optimal statistics discussed in Section 2 (see App. B for details).
[…]
[prior-dependence]
Figure 3: Optimal summaries depend on the prior. Panel (a) shows the parameters of a piecewise likelihood with qualitatively different behaviour on either side of the transition at $\theta=0$. Panel (c) shows two priors with the bulk of their mass on either side of the transition. Panels (b) and (d) show the relationship between the parameter and the sample mean $\bar{y}$ and log variance $\log\operatorname{var}y$, respectively, as a scatter plot. Mutual information estimates highlight that the optimal choice of summary depends on the prior: The $\bar{y}$ and $\log\operatorname{var}y$ summaries are informative for the priors centred at $+1$ and $-1$, respectively.
@@
which is continuous at the transition, as illustrated in panel (a) of Fig. 3. We consider two different normal priors with common standard deviation of $0.25$ centred at ±$\pm 1$, as shown in panel (c). For the purpose of this example, we may choose one summary from the candidate set comprising the sample mean $\bar{y}$ and the natural logarithm of the sample variance $\log\operatorname{var}y$, i.e. we restrict the space of compression functions $\mathcal{T}$ [^2] We restrict $\mathcal{T}$ to simplify the example. Together, the two summaries are sufficient. A continuous mixture of the “left” and “right” part of the likelihood with logistic mixture weight would yield similar results but does not have sufficient statistics.. Intuitively, the latter is informative for the “left” region of the parameter space and the former for the “right”. This intuition is confirmed by simulation: We consider $m=10^{5}$ independent samples from each prior and draw $n=100$ observations from the likelihood in Eq. 11. The relationship between the parameter $\theta$ and sample mean as well as log sample variance are shown in panels (b) and (d), respectively. For quantitative comparison, we also estimate the MI for all pairs of priors and summaries using a nearest-neighbour entropy estimator (Singh et al., 2003). On the one hand, the log sample variance ($\hat{I}=0.70$) is the optimal summary for the left prior because the sample mean provides little information ($\hat{I}=0.02$). On the other hand, the sample mean is highly informative for the right prior ($\hat{I}=0.98$) whereas the log sample variance is not informative ($\hat{I}=0.00$). As Bayesians, we cannot escape the prior, and the optimal summaries depend on it.

### v4

@@ 4.1 Approximate sufficiency
Joyce2008 cast the task of selecting summaries as a sequence of hypothesis tests to select a subset of candidate summaries. Specifically, they considered
$\log R_{k}\left(\theta\right)=\log\tilde{f}\left(\theta\mid t_{k},\ldots,t_{1}\right)-\log\tilde{f}\left(\theta\mid t_{k-1},\ldots,t_{1}\right),$
@@
where $\tilde{f}\left(\theta\mid t_{k-1},\ldots,t_{1}\right)$ is the ABC posterior given k1k-1 summaries already selected and $\tilde{f}\left(\theta\mid t_{k},\ldots,t_{1}\right)$ is the posterior resulting from including an additional statistic $t_{k}$. Intuitively, if the error score $\delta_{k}=\max_{\theta}\operatorname{abs}\left(\log R_{k}\left(\theta\right)\right)$ is zero, i.e. the two posteriors are identical, the $k^{\mathrm{th}}$ statistic does not capture additional information and can be ignored. If $\delta_{k}$ differs significantly from zero, we reject the null hypothesis that $\tilde{f}\left(\theta\mid t_{k},\ldots,t_{1}\right)$ and $\tilde{f}\left(\theta\mid t_{k-1},\ldots,t_{1}\right)$ are the same distribution and include $t_{k}$. They consider a set of k1k-1 summaries to be “approximately sufficient” if $\delta_{k}$ does not significantly differ from zero for any additional summary statistic.
This iterative process cannot minimize a loss functional of the form of Eq. 2 globally. Yet it approximately minimizes a loss functional that assigns all weight to the observed data $y$ and uses the maximum log density ratio to distinguish between true and summary posteriors as the instance-level loss functional, i.e.,
$q(z)$ $=\delta\left(z-y\right)$
$\ell$ $=\max_{\theta}\operatorname{abs}\left(\log f\left(\theta\mid z\right)-\log f\left(\theta\mid t(z)\right)\right),$
where $\delta$ denotes the Dirac delta function.
@@
Importantly, the error score $\delta_{k}=\max_{\theta}\operatorname{abs}\left(\log R_{k}\right)$ assigns equal importance to all subsets of the parameter space, even regions we know to be irrelevant. For example, suppose that the posterior given the currently selected k1k-1 summaries is normal with variance $\sigma_{k-1}^{2}$, and the posterior after adding the $k^{\mathrm{th}}$ summary is identical except for a different variance $\sigma_{k}^{2}$. Even if $\sigma_{k}$ and $\sigma_{k-1}$ differ by an infinitesimal amount, $\delta_{k}$ is unbounded because
@@
$\delta_{k}=\frac{1}{2}\max_{\theta}\operatorname{abs}\left(\log\left(\frac{\sigma_{k}^{2}}{\sigma_{k-1}^{2}}\right)+\left(\frac{\sigma_{k-1}^{2}-\sigma_{k}^{2}}{\sigma_{k}^{2}\sigma_{k-1}^{2}}\right)\theta^{2}\right)=\infty.$
@@
The error score is dominated by regions of the parameter space that have virtually no posterior mass. The expected value $\mathbb{E}_{\theta\sim f\left(\theta\mid t_{k},\ldots,t_{1}\right)}\left[\log R_{k}\left(\theta\right)\right]$ instead weights discrepancies between the two distributions by the posterior mass. This quantity is in fact the KL divergence considered by Barnes2012 (see Section 4.5 for details).
@@
The notion of “approximate” sufficiency is necessarily a statement about limited computational resources: If we had unlimited resources, only candidate statistics that are uninformative or redundant would be excluded. This observation applies to any subset selection algorithm, such as minimizing posterior entropy [Nunes2010] in Section 4.2, regression-based subset selection methods [Blum2010, Blum2013] in Section 4.4, or maximizing MI [Barnes2012] in Section 4.5.
@@ 4.2 Minimizing the conditional posterior entropy
Nunes2010 proposed choosing a subset of summaries $t$ by minimizing the conditional posterior entropy (CPE) $H\left\{f\left(\theta\mid t\left(y\right)\right)\right\}$ given data $y$. They ran rejection ABC for different subsets of summaries and evaluated the CPE using a nearest-neighbor estimator [Singh2003]. The proposal is appealing because low-entropy posteriors give precise parameter estimates.
@@
Figure 2: Extracting summaries can be non-trivial even for toy models. Panel (a) shows the difference between posterior and prior entropy for a model with zero-mean normal likelihood and conjugate gamma prior for the precision $\theta$ (inverse variance). For a subset of the prior and data space, minimizing the posterior entropy discards the second moment $t$, a sufficient statistic. Panel (b) shows the bimodal posterior for the example point in (a) that arises when the precision of the likelihood is $\operatorname{abs}\left(\theta\right)$ (see Section 4.4). The posterior mean is zero and not informative of the parameter. The vertical dashed line represents the maximum likelihood estimate $\widehat{\operatorname{abs}\left(\theta\right)}$ of the precision $\operatorname{abs}\left(\theta\right)$.
@@
where $t=n^{-1}\sum_{i=1}^{n}y_{i}^{2}$ is the second moment, a sufficient statistic. For example, if $a=1.5$ and $t=0.3$, the prior entropy is $1.36$ and the CPE is $1.47$. Minimizing the CPE would discard the sufficient statistic $t$ such that the posterior is equal to the prior: We have not learned anything from the data. Panel (a) of Fig. 2 shows the entropy gain $\Delta=H\left\{f\left(\theta\mid y\right)\right\}-H\left\{\pi\left(\theta\right)\right\}$ in light of the data for different priors and sample variances. Indeed, generating $10^{5}$ samples from the prior predictive distribution with $a=1.5$, we find that $30\%$ of samples lead to a CPE increase. Interestingly, this situation is more likely to arise when the “surprise” [Itti2009] is large, and we should substantially update our beliefs in light of the data. In contrast, the EPE $\mathcal{H}=0.87$ given $t$ is smaller than the prior entropy, and minimizing it would select $t$ as a useful summary. Monte Carlo standard errors of the EPE and proportion of entropy increases are smaller than the reported significant digits.
The instance-level loss functional, the entropy of the summary posterior, is not a discrepancy measure between the true and summary posteriors, and Nunes2010 also considered a two-stage method: First they used the above approach to select candidate summaries and identify simulated datasets close to the observed data. Second, they drew posterior samples for each identified dataset and evaluated the root mean integrated squared error (RMISE) of posterior samples for each subset of summaries. This is possible because the parameters of simulated datasets are known. The summaries with the lowest RMISE were then selected. We do not consider this two-stage approach further here because of its computational burden and because posterior mean estimation methods optimize a similar objective, as discussed in Section 4.4.
@@ 4.3 Maximizing the Fisher information
Even when the likelihood is tractable, compressing the data $y$ to summaries $t$ has computational benefits. Heavens2000 developed an optimal linear compression scheme for Gaussian likelihoods in the sense that the Fisher information is preserved. Information-maximizing neural networks [Charnock2018] seek to maximize the determinant of the Fisher information matrix when linear compression is not sufficient, and methods to maximize the Fisher information for non-Gaussian likelihoods have recently been developed [Alsing2018, Fluri2021]. Fisher information methods are fundamentally likelihood-based and do not fit into the loss functional framework of Eq. 2. However, we can establish a connection to minimizing the EPE in the large-sample limit.
We consider the large-sample limit $n\rightarrow\infty$ of $n$ i.i.d. observations $z=\left(z_{1},\ldots,z_{n}\right)$ and summaries of the form $t\left(z\right)=n^{-1}\sum_{i=1}^{n}h\left(z_{i}\right)$ where $h$ is a potentially nonlinear function. This restriction preserves the i.i.d. structure required for the Bernstein–von Mises theorem and is consistent with the observation that summaries often have well-behaved likelihoods when they are means of i.i.d. data [Alsing2018]. According to the Bernstein–von Mises theorem, the posterior approaches a multivariate normal distribution under certain regularity conditions [Vaart1998]. Specifically,
@@
is the Fisher information of the summaries evaluated at $\theta_{0}$ [Bishop2006, Ch. 6]. The limiting entropy of the posterior can thus be readily evaluated and is
@@
We argue that minimizing the EPE is more appealing than maximizing the Fisher information for three reasons. First, it can incorporate prior information in the small-$n$ regime to yield the most faithful posterior approximation. Second, it does not require the choice of a fiducial value of $\theta$ at which to evaluate the Fisher information. Finally, when the likelihood is not available, we need to approximate it to evaluate the Fisher information. For example, Charnock2018 assume that the likelihood of the learned summaries can be approximated by a Gaussian, and Alsing2018 argue that candidate summaries often have a Gaussian likelihood if they are the mean of i.i.d. data.
@@ 4.4 Minimizing the Bayes risk
Fearnhead2012 proposed the posterior mean of the parameters as summaries. Of course, the posterior mean is not known, but we can estimate it by minimizing the quadratic loss
@@
where $t_{\beta}(z)$ is a predictor of $\theta$ parameterized by $\beta$, $A$ is a positive-definite matrix, and  denotes the transpose. The approach fits into the loss functional framework of Eq. 2 with $q(z)=p\left(z\right)$ (the prior predictive distribution) and instance-level loss functional
$\ell=\int\mathrm{d}z\,f\left(\theta\mid z\right)\left(\theta-t_{\beta}(z)\right)^{\intercal}A\left(\theta-t_{\beta}(z)\right),$
where $t$ is constrained to be the posterior mean. Fearnhead2012 considered linear predictors, but neural networks [Jiang2017] and boosted regression [Aeschbacher2012] have also been proposed. In practice, the parameters $\beta$ are learned by minimizing a Monte Carlo estimate of Eq. 10 akin to Eq. 5. Using the estimated posterior mean $t_{\beta}\left(\cdot\right)$ as summaries implicitly chooses as many summaries as there are parameters.
Considering again the large-sample limit, the quadratic loss becomes (adapted from Theorem 3 of Fearnhead2012)
@@
The above argument crucially depends on the assumptions of the Berstein–von Mises theorem holding. In particular, the model needs to be identifiable such that different values of the parameters $\theta$ are distinguishable in the $n\rightarrow\infty$ limit [Vaart1998]. We consider a variant of the toy model presented in Section 4.2 that is not identifiable and discuss the impact on learning summaries. In particular, we use the absolute value $\operatorname{abs}\left(\theta\right)$ of a parameter $\theta$ as the precision such that the conditional distributions are
$\operatorname{abs}\left(\theta\right)\mid a,b$ $\sim\mathsf{Gamma}\left(a,b\right)$
$y_{i}$ $\sim\mathsf{Normal}\left(0,\operatorname{abs}\left(\theta\right)^{-1}\right).$
The real-valued $\theta$ is distributed as a mixture of a gamma distribution and its reflection about the origin under the prior. The closed-form posterior is
$\operatorname{abs}\left(\theta\right)\mid y,a,b\sim\mathsf{Gamma}\left(a+\frac{n}{2},b+\frac{nt}{2}\right),$
@@
This example may seem contrived, but multimodal posteriors that render the posterior mean uninformative are not uncommon. For example, mixture models are invariant under label permutation [Stephens2000], and latent-space models of networks [Hoff2002] as well as latent factor models for Bayesian PCA [Nirwan2019] are invariant under rotations. The limitation of the Bayes risk approach arises because the instance-level loss functional measures concentration around a point rather than comparing full posterior distributions. Using information theoretic approaches ensures we stay focused on the task at hand: Approximating the true posterior.
The relationship between parameters and data can be complex, and regression approaches, especially linear regression, may not be able to capture the relationship globally. Local relationships in regions of high posterior mass can be learned using pilot runs [Fearnhead2012] or weighting samples [Blum2010]. Local regression methods have also been adapted for subset selection: A model is fit to predict parameters from candidate summaries, and a candidate is selected if it increases a metric such as the Bayesian evidence [Blum2010], Akaike information criterion, or Bayesian information criterion [Blum2013].
@@ 4.5 Maximizing the mutual information
Barnes2012 proposed choosing summaries from a pool of candidates that maximize the MI $I\left\{\theta,t\right\}$ between parameters $\theta$ and the statistics $t$. Assuming that the candidate set includes sufficient statistics $t_{\text{suff}}$ such that
@@
for all possible $y$, they constructed a set of summaries sequentially. At the $k^{\mathrm{th}}$ step, they included the summary that maximizes the surprise given the k1k-1 statistics that have already been selected. The approach is similar to the approximate sufficiency method reviewed in Section 4.1, but candidates are prioritized by their surprise at each stage. Together, the steps select the summaries that maximize the surprise $D_{\mathrm{KL}}\left(f\left(\theta\mid t\right)\;\|\;\pi\left(\theta\right)\right)$ for the observed data. Like Joyce2008 and Nunes2010, this approach considers only the observed dataset with $q(z)=\delta(z-y)$ in the loss functional framework of Eq. 2. Consequently, it maximizes the conditional surprise $\ell=D_{\mathrm{KL}}\left(f\left(\theta\mid t(y)\right)\;\|\;\pi\left(\theta\right)\right)$ rather than the MI, which is the expected surprise under the prior predictive distribution.
However, recall from Eqs. 6 and 7 that the MI is equal to the expected surprise under the generative model. In general, maximizing the surprise for a particular observed dataset is thus not equivalent to maximizing the MI. The approach may select different summaries if the candidate set does not include sufficient statistics.
Similarly, Chen2021MutualInfo sought to maximize the MI using a neural network. They suggested that “$t(z)$ is a sufficient statistic for $g\left(z\mid\theta\right)$ if and only if” (p. 2) it maximizes the MI and “that the sufficiency of the learned statistics is insensitive to the choice of $\pi\left(\theta\right)$” (p. 4) such that “[their approach] is globally sufficient for all $\theta$” (p. 6) [^1] We have adapted notation in quotations for consistency with this manuscript.. As we shall illustrate with a toy model, these propositions do not hold in general because of the difference between sufficient and optimal statistics discussed in Section 2 (see App. B for details).
Figure 3: Optimal summaries depend on the prior. Panel (a) shows the parameters of a piecewise likelihood with qualitatively different behaviour on either side of the transition at $\theta=0$. Panel (c) shows two priors with the bulk of their mass on either side of the transition. Panels (b) and (d) show the relationship between the parameter and the sample mean $\bar{y}$ and log variance $\log\operatorname{var}y$, respectively, as a scatter plot. Mutual information estimates highlight that the optimal choice of summary depends on the prior: The $\bar{y}$ and $\log\operatorname{var}y$ summaries are informative for the priors centred at +1+1 and 1-1, respectively.
@@
which is continuous at the transition, as illustrated in panel (a) of Fig. 3. We consider two different normal priors with common standard deviation of $0.25$ centred at $\pm 1$, as shown in panel (c). For the purpose of this example, we may choose one summary from the candidate set comprising the sample mean $\bar{y}$ and the natural logarithm of the sample variance $\log\operatorname{var}y$, i.e. we restrict the space of compression functions $\mathcal{T}$ [^2] We restrict $\mathcal{T}$ to simplify the example. Together, the two summaries are sufficient. A continuous mixture of the “left” and “right” part of the likelihood with logistic mixture weight would yield similar results but does not have sufficient statistics.. Intuitively, the latter is informative for the “left” region of the parameter space and the former for the “right”. This intuition is confirmed by simulation: We consider $m=10^{5}$ independent samples from each prior and draw $n=100$ observations from the likelihood in Eq. 12. The relationship between the parameter $\theta$ and sample mean as well as log sample variance are shown in panels (b) and (d), respectively. For quantitative comparison, we also estimate the MI for all pairs of priors and summaries using a nearest-neighbor entropy estimator [Singh2003]. On the one hand, the log sample variance ($\hat{I}=0.70$) is the optimal summary for the left prior because the sample mean provides little information ($\hat{I}=0.02$). On the other hand, the sample mean is highly informative for the right prior ($\hat{I}=0.98$) whereas the log sample variance is not informative ($\hat{I}=0.00$). As Bayesians, we cannot escape the prior, and the optimal summaries depend on it.